OpenStudy (anonymous):

Solve: 5(x + 7)2 = 25

OpenStudy (kymber):

You mean 5(x + 7)² = 25 ?

jimthompson5910 (jim_thompson5910):

5(x + 7)^2 = 25 (x + 7)^2 = 25/5 (x + 7)^2 = 5 x+7 = sqrt(5) or x+7 = -sqrt(5) x = -7+sqrt(5) or x = -7-sqrt(5)